# PCG-Informed Neural Solvers for High-Resolution Homogenization of Periodic Microstructures

### v1

@@ 2.4. PCG-Informed learning: joint optimization with embedded PCG refinement
iteration $\bm{\delta_{Max}}$ $\bm{\delta_{Min}}$ $\bm{\delta_{Mean}}$ $\bm{||r||_{Mean}}$ Time (ms)
@@
8 $0.91\%$ $0.15\%$ $0.51\%$ $3.38\times 10^{-3}$ 158.27
[…]
Joint Training with PCG to Suppress Spectral Bias.
Accurately capturing displacement fields in microstructures with complex topology presents a significant challenge for neural solvers. This difficulty arises from the spectral bias of neural networks—their inherent tendency to prioritize smooth, low-frequency features while underrepresenting complex, high-frequency spatial variations. In topologically intricate structures such as Truss lattices, where the displacement fields exhibit highly irregular and spatially rich patterns due to the geometric complexity, conventional neural models often fail to produce stable or physically consistent predictions.
To mitigate this, CGiNS integrates a lightweight PCG solver directly into the decoder and trains the entire architecture in a solver-in-the-loop fashion. Unlike post-training numerical corrections, this joint optimization ensures that the network produces intermediate predictions already compatible with the iterative dynamics of PCG. As a result, only a few refinement steps are required to suppress residuals and recover physically meaningful solutions.
@@
This demonstrates that CGiNS can effectively suppress spectral bias through minimal iterative correction, enabling accurate, stable, and efficient predictions across a wide range of structural complexities.
@@
This form reveals that the total projection energy is a weighted sum of the error components, where each component is scaled by the corresponding eigenvalue $\lambda_{i}$. High-energy components (associated with large $\lambda_{i}$) dominate the error and are harder to eliminate in early PCG iterations, which tend to smooth out low-frequency (small $\lambda$) components more efficiently. Therefore, a good network initialization should minimize $\alpha_{i}$ in high-$\lambda$ directions, aligning the output with the low-frequency modes of $\tilde{K}$.
@@ 3. Discussion
This work introduces CGiNS, a hybrid neural-PDE framework for fast and accurate homogenization of periodic microstructures. By embedding physics-based priors—such as periodic boundary conditions and equilibrium constraints—directly into the model architecture, CGiNS achieves high-resolution generalization, low residuals, and strong physical consistency, setting a new standard in learning-based homogenization.
A central component of CGiNS is the integration of a PCG module during training. This embedded solver serves a dual purpose: it refines predicted displacement fields through a small number of iterations and simultaneously trains the network to produce PCG-friendly predictions that are well-aligned with the numerical solution process. As a result, the model achieves fast convergence while preserving the fidelity of elasticity responses, even under complex topologies.
Beyond the PCG component, CGiNS incorporates several architectural modules that, while designed for homogenization, embody principles broadly applicable to other physics-based learning systems. The Peri-mapping data structure enables periodic boundary conditions by allowing convolutions to access features from corresponding nodes across domain boundaries. This direct encoding of cyclic connectivity is broadly applicable to spatial models requiring periodicity, such as lattice-based systems or wave simulations. The soft voxel strategy addresses a key challenge in sparse representations: loss of boundary detail due to limited receptive fields. By adding low-stiffness padding, it preserves geometric detail near interfaces with minimal overhead, offering a general solution for boundary fidelity in sparse learning tasks. The multi-level prediction framework, inspired by multigrid methods, enables efficient learning across scales and improves convergence in high-resolution settings. Its hierarchical design can benefit other PDE solvers with multi-scale structure. Together, these components illustrate a broader design principle: physical consistency, geometric fidelity, and numerical efficiency can be jointly achieved through solver-aware architectural integration.
Despite these advantages, several limitations remain. This work focuses on problems governed by Dirichlet-type boundary conditions. While many physical systems also involve Neumann or mixed boundary conditions, these have not been considered in our current formulation. In addition, although CGiNS demonstrates resolution scalability up to $512^{3}$, this upper bound is primarily due to GPU memory constraints rather than architectural limitations. With further engineering—such as mixed-precision computation, domain decomposition, or memory-efficient modeling—the framework could be extended to even higher resolutions.
Future work will explore extending CGiNS as a general-purpose neural PDE solver, beyond its current application to periodic homogenization. One direction is to support a wider variety of boundary conditions—including Neumann and mixed types. Another avenue is to replace the fixed preconditioner with a learnable, potentially improving convergence robustness. Finally, we aim to generalize the framework to solve a broader class of elliptic and steady-state PDEs, using CGiNS as a foundation for solver design rather than application-specific modeling.
In conclusion, CGiNS shows that integrating solver-aware constraints into the learning pipeline can significantly improve reliability and scalability. Its success encourages deeper integration between learning and simulation in the development of next-generation PDE solvers.
@@ 4.3. Loss function
Formally, let $k(K)$ denote the dimensionality of the stiffness matrix $K$, $D(K)$ the Jacobian preconditioner used in the PCG solver, and $P_{t}(\cdot)$ the result after $t$ iterations of PCG. Let $u$ denote the predicted displacement field and $\bar{u}$ its mean over all nodes. To enforce translation-invariant displacement predictions, we define $\widetilde{u}=u-\bar{u}$. The energy-based loss function at the finest resolution is then formulated as:
@@ Appendix A Homogenization theory
Here, $C^{H}_{ijkl}$ represents the elastic tensor of the base material with Voigt notation. $i,j,k,l\in\{1,2,3\}$, where 1, 2, and 3 represent the $x$, $y$, and $z$ directions, respectively. $\overline{\epsilon}_{pq}$ is the prescribed macroscopic stress field, in this work, we consider the following six strain fields:
@@
$v$ is a virtual displacement field. $C^{b}$ is the elasticity tensor of the microstructure base material, which is isotropic and depend on the base material’s Young’s modulus $E$ and Poisson’s ratio $\nu$. We define $\Theta=\frac{\nu}{1-\nu}$ and $\Phi=\frac{1-2\nu}{2(1-\nu)}$. It can be computed as:
(16) $\begin{split}C^{b}&=\mathcal{M}(E,\nu)\\ &=\frac{E(1-\nu)}{(1+\nu)(1-2\nu)}\begin{bmatrix}1&\Theta&\Theta&0&0&0\\ \Theta&1&\Theta&0&0&0\\ \Theta&\Theta&1&0&0&0\\ 0&0&0&\Phi&0&0\\ 0&0&0&0&\Phi&0\\ 0&0&0&0&0&\Phi\end{bmatrix}.\end{split}$
@@
After discretizing the microstructure, assign degree-of-freedom (DOF) labels. Assemble the global stiffness matrix $K(\Omega,\mathcal{M})$ and global load vector $f(\Omega,\mathcal{M})$ by mapping elemental stiffness matrices $K_{e}$ and loads vector $f_{e}$ through DOF connectivity (local-to-global node numbering). Solve the linear system to obtain the displacement field $u$:
@@ F-RD $4(\cos x\cos y\cos z)-(\cos 2x\cos 2y+\cos 2y\cos 2z+\cos 2x\cos 2z)=c$
Truss microstructures are periodic frameworks composed of interconnected truss units, designed to achieve ultralightweight yet mechanically robust architectures (Liu et al., 2008). Optimized spatial configurations enable remarkable specific strength and stiffness, making trusses a staple in engineering design, lightweight materials, and advanced manufacturing.
@@ B.2. Dataset construction
We perform a t-SNE dimensionality reduction on the $6\times 6$ effective elasticity tensor $C^{H}$ computed from all microstructure samples, projecting the data into a 2D space for visualization. As shown in Fig. 6, each point corresponds to a sample in one of the three datasets: TPMS (blue), PSL (red), or Truss (green). As shown, the three datasets form distinguishable yet partially overlapping clusters, reflecting their differences in topological regularity, geometric complexity, and resulting mechanical behavior.
@@
Table 8. Overview of microstructure datasets with topological and volume fraction variations. This table summarizes the datasets used for training and evaluation. Each dataset represents periodic microstructures with varying topology and volume fraction (VF). For TPMS, topology refers to five distinct implicit surface equations (see Tab.7); for PSL and Truss, it refers to different skeletal graph structures derived from procedural designs or engineering templates. Each dataset includes separate splits for training, test, and validation. Poisson’s ratios $\nu$ are randomly sampled from $[0.2,0.4]$, and for benchmarking against baseline methods, material properties are fixed at $E=1$, $\nu=0.3$. The volume fraction for all samples ranges from $10\%$ to $40\%$.
@@ Appendix C Additional validations
[…]
C.1. Performance of prediction accuracy
@@ C.3. Comparison with larger PH-Net
C.4. Inference speed of neural baselines
Table 12. Inference time of neural solvers on resolution $64$ TPMS structures. Although baseline methods exhibit low latency, they fail to provide physically reliable outputs (see main results). CGiNS offers a balanced trade-off between speed and solution accuracy.

### v2

@@ 2.4. PCG-Informed learning: joint optimization with embedded PCG refinement
[…]
@@
8 $0.91\%$ $\% 0.510.51\%$ $3.38\times 10^{-3}$ 158.27
10 $\bm{0.70\%}$ $\bm{0.11}\%$ $\bm{0.50\%}$ $\bm{3.34\times 10^{-3}}$ 158.91
Joint Training with PCG
Accurately capturing displacement fields in microstructures with complex topology presents a significant challenge for neural solvers. To address this, CGiNS integrates a lightweight PCG solver directly into the decoder and trains the entire architecture in a solver-in-the-loop fashion. Unlike post-training numerical corrections, this joint optimization ensures that the network produces intermediate predictions already compatible with the PCG. As a result, only a few refinement steps are required to suppress residuals and recover physically meaningful solutions.
@@
In the preceding sections we leveraged embedded PCG primarily as a lightweight refinement to suppress the network’s spectral bias. Here we adopt a complementary view of their interplay: rather than merely separating high and low frequency errors, the neural network can serve PCG by providing a high-quality initial guess that is already well aligned with the geometry of the preconditioned system. When $u_{0}$ lies mostly in favorable eigendirections, PCG needs only a few iterations to remove a small, well-conditioned residue. To make this notion precise, we quantify initialization quality with a simple projection energy metric.
@@
This form reveals that the total projection energy is a weighted sum of the error components, where each component is scaled by the corresponding eigenvalue $\lambda_{i}$. In PCG, the convergence rate depends strongly on the spectral distribution of the preconditioned system. Error components aligned with eigen-directions of higher energy (large $\lambda$) are typically harder to eliminate, meaning that if the initial solution carries large projection weights in these directions, convergence will be slow. Conversely, components with lower energy (small $\lambda$) are reduced more quickly, so a good initialization is one that minimizes the error energy in the high-$\lambda$ directions. In this sense, the projection energy provides a direct measure of initialization quality and its impact on subsequent PCG efficiency.
@@ 2.5. Ablation study
2.6. Multiphase homogenization case
Table 7. Multiphase homogenization with CGiNS on the Truss dataset. Results for test cases with varying numbers of phases (2, 3, 5, and INF, where INF denotes fully random phase assignment). We recorded the relative error of the computed solution with respect to the reference solution ($\delta_{Mean}$), and the residual norm of the linear system ($\|r\|_{Mean}$).
Metrics Number of Phases
2 3 5 INF
$\delta_{Mean}$ $0.33\%$ $0.65\%$ $0.90\%$ $1.01\%$
$\|r\|_{Mean}$ $2.11\times 10^{-3}$ $5.52\times 10^{-3}$ $4.83\times 10^{-3}$ $6.09\times 10^{-3}$
In line with practical manufacturing, our main experiments emphasize single-phase lattices, which dominate current fabrication tasks, have mature design–for–manufacture pipelines, and are standard building blocks for lightweighting, energy absorption, and biomedical implants; periodic unit-cell single-phase lattices of this kind are ubiquitous in industrial workflows.
Importantly, CGiNS encodes material parameters at each voxel and thus naturally extends to multiphase homogenization without modifying the overall pipeline or requiring retraining. To illustrate this capability, we constructed an dataset of 10,000 Truss microstructures in which each voxel can take a distinct material: 2-phase, 3-phase, 5-phase, and an “INF” case with randomly phased assignments (2,500 samples per category). Materials are non-dimensionalized with $E=1$ and $\nu\in\left[0.2,0.4\right]$. We conducted additional tests on benchmark datasets, using CGiNS pretrained on single-phase materials (Table 2’s pretrained model).
As summarized in Table 7, mean residual norms for the linear solves are on the order of $10^{-3}$, while the mean relative error of the elasticity tensor remains below $1\%$ for 2-phase ($0.33\%$), 3-phase ($0.65\%$), and 5-phase ($0.90\%$) cases; the INF setting yields $1.01\%$, a slight increase attributable to strong material jumps across neighboring voxels. These results indicate that CGiNS maintains high accuracy for multiphase homogenization and is readily extensible beyond single-phase microstructures.
@@ 3. Discussion
This work introduces CGiNS, which tightly couples geometry-aware sparse convolutions, strict periodic encoding, and a lightweight PCG refinement to jointly achieve accuracy, stability, and efficiency for high-resolution periodic microstructure homogenization. Unlike end-to-end regressors or approaches with weak physical priors, CGiNS encodes periodic boundaries and the global displacement constraint explicitly in the network, and then aligns the prediction with the dynamics of a numerical solver via a few PCG steps during inference. This coupling yields two immediate benefits: (i) errors of the effective elasticity tensor remain below $1\%$ with residuals on the order of $1\times 10^{-3}$ across complex topologies; and (ii) end-to-end speedups of $2\times$ to $10\times$ over GPU multigrid solvers while scaling to $512^{3}$resolutions, where hybrid methods often suffer from memory or convergence issues.
Across three representative datasets (TPMS, PSL, and Truss), CGiNS consistently reduces effective-tensor error and improves numerical stability compared with 3D-CNN, Label-Free, PH-Net, and other baselines. The gains are most pronounced on geometrically irregular cases (PSL, Truss) with strong high-frequency content in the displacement field, where baselines frequently exceed $10\%$ error or become unstable, whereas CGiNS maintains $\textless 1\%$ error and lower residuals. These results indicate that a tightly coupled “Neural + Numerical” pipeline mitigates spectral bias and preserves physical consistency.
Across resolutions from $64^{3}$ to $512^{3}$, CGiNS achieves $2\times$ to $10\times$ end-to-end speedups over GPU multigrid solvers (AmgX/Homo3D) at matched accuracy thresholds, and at $512^{3}$ the overall wall-clock improvement reaches $8.6\times$ to $11.9\times$ on TPMS/PSL/Truss. Notably, once the homogenized tensor error is around $1\%$, further driving the linear-system residual lower with classical solvers yields negligible gains in $C^{H}$, which explains CGiNS ’s favorable accuracy–time trade-off: a few physically consistent PCG steps suffice to reach industrial-grade accuracy. At the same time, results with learned numerical hybrids—e.g., learning the AMG P-operator with GNNs—reveal that a plain ”network ++ solver” recipe is not enough in complex settings. While such methods can accelerate regular geometries (e.g., TPMS), they degrade in success rate and efficiency on irregular, high-contrast cases (PSL, Truss). The core issue is a lack of geometry-aware inductive bias: without explicitly encoding periodicity, boundary handling, and global constraints, the learned transfer/relaxation operators become fragile under geometric variability. CGiNS addresses this gap by embedding geometry and physics directly into the learning system—via geometry-aware sparse convolutions, strict periodic encoding, and a global displacement constraint—then aligning the prediction with a lightweight in-the-loop PCG. This combination preserves multigrid-like cross-scale structure while remaining robust on complex topologies, delivering both the stability that learned hybrids struggle to maintain and the speedups that classical solvers alone cannot match at comparable accuracy.
Ablations clarify the role and limits of each component: (1) PCG-in-the-loop. Injecting a small number of PCG steps after the network prediction markedly reduces residuals and restores physically interpretable solutions; benefits plateau around  10 steps. Removing PCG causes sharp error increases on complex topologies (Truss), showing it is essential for convergence and stability rather than a cosmetic refinement. (2) Multi-level prediction. Increasing from one to two levels significantly lowers error and residual; a third level shows diminishing returns, underscoring the need for cross-scale features for high-fidelity displacement fields. (3) Global displacement constraint. Enforcing a zero-mean displacement in lieu of heuristic Dirichlet anchors improves consistency and transferability on complex meshes, avoiding anchor ambiguity. (4) Soft-voxel boundary buffering and peri-mapping for periodicity improve boundary-geometry fidelity under sparse representations and preserve periodic consistency; ablating either degrades errors and residuals in a systematic way.
Regarding material generalization and multi-phase extensions, CGiNS encodes isotropic matrix parameters explicitly as inputs, enabling stable predictions across material settings. Even for 2/3/5-phase and voxel-level random phase assignments (INF), errors of the effective stiffness remain around $1\%$. This suggests that a self-supervised energy-minimization loss combined with PCG-in-the-loop training preserves physical consistency and numerical robustness across the coupled geometry–material space.
Limitations and outlook.
This work targets linear elasticity with periodic boundaries and a global displacement constraint (Dirichlet-equivalent). More general Neumann/mixed boundaries, nonlinear materials, and contact are not yet covered. The current $512^{3}$ ceiling is mainly due to memory and implementation limits; future improvements include mixed precision, domain decomposition, and more memory-efficient operators. A key direction is cross-resolution prediction: train at a single “base” resolution and deploy zero or few-shot across other resolutions using scale-conditioned encodings plus multigrid-consistent restriction and prolongation to preserve invariants. Another avenue is replacing fixed preconditioners with learnable yet stability-controlled modules, and extending the framework to broader PDEs. The overarching lesson remains: maintain a physically consistent feedback between the network and the numerical iteration to reach industrial-grade accuracy with minimal extra solves.
@@ 4.3. Loss function
Formally, let $k(K)$ denote the dimensionality of the stiffness matrix $K$, $D(K)$ the Jacobi preconditioner used in the PCG solver, and $P_{t}(\cdot)$ the result after $t$ iterations of PCG. Let $u$ denote the predicted displacement field and $\bar{u}$ its mean over all nodes. To enforce translation-invariant displacement predictions, we define $\widetilde{u}=u-\bar{u}$. The energy-based loss function at the finest resolution is then formulated as:
@@ Appendix A Homogenization theory
Here, $C^{H}_{ijkl}$ represents the elastic tensor of the lattice microstructure with Voigt notation. $i,j,k,l\in\{1,2,3\}$, where 1, 2, and 3 represent the $x$, $y$, and $z$ directions, respectively. $\overline{\epsilon}_{pq}$ is the prescribed macroscopic strain field, in this work, we consider the following six strain fields:
@@
$v$ is a virtual displacement field. $C^{b}$ is the elasticity tensor of the microstructure base material, which is isotropic and depends on the base material’s Young’s modulus $E$ and Poisson’s ratio $\nu$. We define $\Theta=\frac{\nu}{1-\nu}$ and $\Phi=\frac{1-2\nu}{2(1-\nu)}$. It can be computed as:
[…]
@@
After discretizing the microstructure, we assign degree-of-freedom (DOF) labels. Assemble the global stiffness matrix $K(\Omega,\mathcal{M})$ and global load vector $f(\Omega,\mathcal{M})$ by mapping elemental stiffness matrices $K_{e}$ and load vector $f_{e}$ through DOF connectivity (local-to-global node numbering). Solve the linear system to obtain the displacement field $u$:
@@ F-RD $4(\cos x\cos y\cos z)-(\cos 2x\cos 2y+\cos 2y\cos 2z+\cos 2x\cos 2z)=c$
Truss microstructures are periodic frameworks composed of interconnected truss units, designed to achieve ultra-lightweight yet mechanically robust architectures (Liu et al., 2008). Optimized spatial configurations enable remarkable specific strength and stiffness, making trusses a staple in engineering design, lightweight materials, and advanced manufacturing.
@@ B.2. Dataset construction
We performed a t-SNE dimensionality reduction on the $6\times 6$ effective elasticity tensor $C^{H}$ computed from all microstructure samples, projecting the data into a 2D space for visualization. As shown in Fig. 6, each point corresponds to a sample in one of the three datasets: TPMS (blue), PSL (red), or Truss (green). As shown, the three datasets form distinguishable yet partially overlapping clusters, reflecting their differences in topological regularity, geometric complexity, and resulting mechanical behavior.
@@
Table 9. Overview of microstructure datasets with topological and volume fraction variations. This table summarizes the datasets used for training and evaluation. Each dataset represents periodic microstructures with varying topology and volume fraction (VF). For TPMS, topology refers to five distinct implicit surface equations (see Table 8); for PSL and Truss, it refers to different skeletal graph structures derived from procedural designs or engineering templates. Each dataset includes separate splits for training, test, and validation. Poisson’s ratios $\nu$ are randomly sampled from $[0.2,0.4]$, and for benchmarking against baseline methods, material properties are fixed at $E=1$, $\nu=0.3$. The volume fraction for all samples ranges from $10\%$ to $40\%$.
@@ Appendix C Additional validations
TPMS $10\%-20\%$ $17.57\%$ $6.08\%$ $10.22\%$ $44.81\%$ $16.11\%$ $22.68\%$ $2.34\%$ $0.88\%$ $1.43\%$ $\bm{0.18\%}$ $\bm{0.00\%}$ $\bm{0.06\%}$
$20\%-30\%$ $11.98\%$ $5.22\%$ $9.04\%$ $32.49\%$ $14.31\%$ $32.49\%$ $2.62\%$ $0.86\%$ $1.49\%$ $\bm{0.09\%}$ $\bm{0.02\%}$ $\bm{0.04\%}$
$30\%-40\%$ $11.51\%$ $2.10\%$ $6.36\%$ $19.80\%$ $10.86\%$ $15.36\%$ $2.53\%$ $0.85\%$ $1.44\%$ $\bm{0.12\%}$ $\bm{0.02\%}$ $\bm{0.04\%}$
PSL $10\%-20\%$ $106.64\%$ $23.76\%$ $51.87\%$ $226.26\%$ $46.86\%$ $100.81\%$ $473.37\%$ $4.92\%$ $30.08\%$ $\bm{0.64\%}$ $\bm{0.22\%}$ $\bm{0.32\%}$
$20\%-30\%$ $111.99\%$ $12.20\%$ $41.47\%$ $239.98\%$ $43.24\%$ $92.86\%$ $308.04\%$ $5.52\%$ $23.60\%$ $\bm{0.55\%}$ $\bm{0.19\%}$ $\bm{0.32\%}$
$30\%-40\%$ $73.73\%$ $8.57\%$ $27.13\%$ $155.86\%$ $30.54\%$ $68.85\%$ $188.32\%$ $4.12\%$ $15.90\%$ $\bm{0.38\%}$ $\bm{0.14\%}$ $\bm{0.29\%}$
Truss $10\%-20\%$ $197.54\%$ $16.02\%$ $66.94\%$ $836.25\%$ $54.58\%$ $278.99\%$ $104.04\%$ $4.90\%$ $65.67\%$ $\bm{0.84\%}$ $\bm{0.32\%}$ $\bm{0.64\%}$
$20\%-30\%$ $142.22\%$ $9.86\%$ $43.94\%$ $463.34\%$ $40.61\%$ $158.90\%$ $357.14\%$ $3.25\%$ $36.31\%$ $\bm{0.73\%}$ $\bm{0.19\%}$ $\bm{0.51\%}$
$30\%-40\%$ $98.90\%$ $7.66\%$ $30.52\%$ $249.64\%$ $33.17\%$ $94.53\%$ $329.14\%$ $2.58\%$ $18.43\%$ $\bm{0.43\%}$ $\bm{0.10\%}$ $\bm{0.38\%}$
C.1. Prediction accuracy
@@ C.3. Comparison with larger PH-Net
C.4. Inference speed of neural homogenization baselines
Table 13. Inference time of neural solvers on resolution $64$ Truss structures. Although baseline methods exhibit low latency, they fail to provide physically reliable outputs (see main results). CGiNS offers a balanced trade-off between speed and solution accuracy.